OpenStudy (hesan):

If x is a positive number, what is the least value of x^2-2x+3?

OpenStudy (anonymous):

fo minima,double derivative of function wrt to x is negative

OpenStudy (anonymous):

sorry positive

OpenStudy (hesan):

what will be the value then?

OpenStudy (anonymous):

2x-2=0 2x=2 x=1 for minima, f(1)=1-2+3 =4-2 =2 in this question,i am assuming f(x)=x^2-2x+3

OpenStudy (hesan):

But the answer isn't 2

OpenStudy (anonymous):

least value is 1

OpenStudy (hesan):

No it is't 1 either...

OpenStudy (anonymous):

no sorry i am confused wait

OpenStudy (anonymous):

find the first derivative , make it equal to zero u will get x=1 then find the second derivative by putting the value of x=1 in it if value of second derivative is positive then we will say that the given function has a minimum value at x=1 to find least value put x=1 in original equation u will get the answer as 2

ganeshie8 (ganeshie8):

is it 3 ?

OpenStudy (hesan):

Nup

OpenStudy (anonymous):

2 has to be the answer

ganeshie8 (ganeshie8):

this Q from calculus, or quadratic equations ?

OpenStudy (hesan):

Quadratic eq

OpenStudy (hesan):

The answer given is 0 ?

ganeshie8 (ganeshie8):

x coordinate of vertex = -b/2a = 2/2 = 1 so, x is a positive number. y coordinate of vertex is the minimum value. can u find that ?

OpenStudy (hesan):

you mean 0 is the wrong answer, it should be 2?

ganeshie8 (ganeshie8):

@nitz is right i am also getting 2.. ur book is wrong... yeah :)

OpenStudy (hesan):

okey... thnx

ganeshie8 (ganeshie8):

np :)

OpenStudy (anonymous):

the final answer is 2